What is the mass of Co(NO3)2 in 367 mLs of solution that contains only 0.551 moles of Co(NO3)2?
Answer #1

As per your questions given,

Moles of Co(NO3)2 = 0.551 moles

Molar mass of the compound = 59 + 2×(14 + 48) = 183 g/mol

Mass = molar mass× moles

So mass of compound = 0.551 × 183 = 100.83 gram